tanA+tanB=

tanA+tanB=

Grade:12

9 Answers

Gavvala Ganesh
8 years ago
tanA+tanB/1-tanAtanB
SHANMUKESHWAR
8 years ago
tanA+tanB/1-tanAtanB
Raghu Vamshi Hemadri
8 years ago
tanA+tanB/1-tanAtanB
L GOUSE BASHA
8 years ago
tanA+tanB/1-tanAtanB
Nikhil Upadhyay
8 years ago
All the above 4 answers are wrong. The correct Answer is tan A+ tanB=tan(A+B)*(1-tanAtanB).PLZ CONSIDER THIS>
OUTPASS
8 years ago
tanA+tanB/1-tanAtanB.
KARTHIK
8 years ago
tanA+tanB/1-tanAtanB
saibhargav
8 years ago
tanA+tanB/1-tanAtanB
SHAIK HAFEEZUL KAREEM
8 years ago
tanA+tanB=tan(A+B)*(1-tanAtanB)
this is correct answer cone from the sum of angle formulas i.e.,tan(A+B)=tanA+tanB/1-tanAtanB